### v1

# A Review of Community-Centric Power Systems Resilience Assessment and Enhancement Strategies

Abstract
This paper presents a comprehensive review of resilience metrics, covering both engineering-based measures, such as fragility-curve modeling, and data-driven approaches, including triangular and trapezoidal representations. Next, the paper examines the interdependencies between power systems resilience and community resilience, addressing socioeconomic and behavioral dimensions, infrastructure interconnections, and the emerging role of resilience hubs. The review then synthesizes state-of-the-art strategies for enhancing power system resilience, including network hardening, resource allocation, optimal scheduling, and reconfiguration techniques. Special emphasis is placed on the integration of Artificial Intelligence (AI) methods and the techno-legal dimensions of resilient power systems and communities. In particular, the paper contrasts the regulatory landscapes of the European Union and the United States, highlighting key similarities and distinctions. By analyzing methodologies for mitigating the impacts of high-impact, low-probability (HILP) events, the review identifies critical research gaps and outlines promising directions for future investigation.
keywords:
Community resilience, high-impact low-probability events, power systems resilience, resilience evaluation methods, resilience enhancement strategies, resilience metrics.
1 Introduction
The frequency of power outages and their associated losses have been growing in the U.S. over the past 43 years, as illustrated in Figure 1[1]. Aging infrastructure and the rise of extreme events have highlighted the vulnerability of power systems [2]. Storm-related disruptions are estimated to cost the U.S. economy between \$20 billion to \$55 billion annually [3]. Several research studies anticipate the frequency and severity of such extreme events to increase due to climate change [3, 4, 5]. This underscores the critical need to adapt power grids to withstand disasters and mitigate catastrophic impacts. Ensuring the resilience of the energy supply and infrastructure is essential for facilitating economic recovery and fostering sustainable development.
Various definitions, metrics, assessment methods, and improvement strategies for power system resilience have been proposed. However, the absence of standardized definitions and metrics poses a challenge in accurately estimating the resilience of power systems [6]. The U.S. Department of Energy (DOE) defines energy system resilience as “the ability of the grid, buildings, and communities to withstand and rapidly recover from power outages while continuing to provide electricity”. According to the IEEE, resilience is defined as “the ability to withstand and reduce the magnitude and/or duration of disruptive events, which includes the capability to anticipate, adapt to, and/or rapidly recover from such an event.” Community resilience builds upon this concept by extending it to social and infrastructural systems, emphasizing collective preparedness and adaptive capacity which is defined as a community’s ability to reduce and prepare for risks [83]. Berkes et al. [84] proposed two perspectives on community resilience: the social-ecological systems perspective and the psychology and community development perspective. Various socioeconomic and demographic factors that shape community resilience also influence power system resilience [70, 72]. Additionally, power system resilience assessments must consider the interdependency of power systems with other critical infrastructure, as cascading failures can impact multiple systems. Proper parameter tuning is essential for maintaining the stability of integrated networks [81, 82].
Several research studies have extensively analyzed the concepts of resilience and strategies for enhancement, providing a detailed overview encompassing definitions, evaluation methodologies, modeling approaches, and strategies for enhancing power system resilience [9, 10, 11, 12]. Despite these contributions, there is a lack of comprehensive exploration of the resilience triangle and trapezoid frameworks, detailed assessment methodologies, and the integration of modeling techniques with AI and ML for resilience enhancement. Additionally, comprehensive discussions linking policy, regulatory frameworks, and socio-economic factors to resilience enhancement remain limited. This paper addresses these gaps by providing an in-depth analysis of power system resilience evaluation and enhancement methods against extreme weather events, with a primary focus on the state-of-the-art AI and ML techniques.
The structure of the paper is organized as follows. Section 2 explores resilience assessment approaches. Section 3 presents a discussion on the interdependence of power system resilience and community resilience. Section 4 delves into resilience enhancement strategies, then Section 5 presents the application of AI in resilience analysis. Section 6 reviews the techno-legal aspects of resilient power systems and communities. Section 7 identifies research gaps, outlines challenges, and proposes guidance for future research. Finally, Section 8 concludes the paper with a summary of the key findings and a discussion of the identified limitations within current research efforts.
Figure 1: Frequency and associated economic loss of U.S. extreme weather events.
2 Power System Resilience Assessment Practices
Effective implementation of resilient planning and operations depends on a comprehensive evaluation of resilience using standardized metrics. Methods for assessing resilience in power systems have received increasing attention in recent research and can be broadly classified into two main approaches: 1) Detailed engineering (micro-level), and 2) Data-driven (macro-level).
2.1 Detailed Engineering Resilience Assessment
Detailed engineering resilience assessment techniques emphasize analyzing the specific fragility curves of individual components within a system, offering a granular perspective on their vulnerability to extreme events. A fragility curve represents the probability of failure of a system or one of its components under a specific event. In scenarios with minimal uncertainty, the fragility curve closely resembles a step function, transitioning sharply between two states: 0 (non-failure) and 1 (failure) [32]. However, in most real-world applications, fragility curves are modeled using lognormal cumulative distribution functions, which rely on the logarithmic mean and logarithmic standard deviation to provide a more realistic representation by accounting for variations in system behavior under different load conditions. In [34], the fragility curve is described as a lognormal distribution
$f(x)=\frac{l}{\sigma\sqrt{2\pi}}e^{-\frac{1}{2}(\frac{x-\mu}{\sigma})^{2}},$ (1)
where $x$ represents event intensity, $\mu$ and $\sigma$ denote the logarithmic mean and standard deviation, respectively. The slope of the fragility curve is influenced by these two parameters, determining the system’s sensitivity to variations in event intensity. For example, Figure 2 illustrates fragility curves for transmission lines and towers under varying wind speeds (a) and infrastructure failure as a function of peak ground acceleration (PGA) during earthquakes (b). The probability of failure ($P_{f}$) in seismic scenarios can be expressed as:
$P_{f}(PGA)=\alpha\left[\frac{1}{\beta_{tot}}\ln\left(\frac{PGA}{PGA_{mi}}\right)\right],$ (2)
where $P_{f}$ denotes the probability of infrastructure failure, $\alpha$ is the standard cumulative probability function, $PGA_{mi}$ is the median threshold value of the peak ground acceleration, and $\beta_{tot}$ is the total standard deviation [35].
Figure 2: Fragility curve vs extreme event intensity as a function of wind speed (a), and of PGA (b).
Ranjbar et al. [19] analyzed fragility curves for substations, wind farms, and transmission lines during hurricanes. Substation damage probabilities were modeled using log-normal distributions, transmission structure failures with exponential functions, and wind turbine failures with log-logistic distributions based on wind speeds. Four methodologies have been established for constructing fragility curves for power system components: 1) judgmental, based on expert experience, 2) empirical, based on field and laboratory measurements to model the performance of the components [33], 3) analytical, derived from structural simulation models, and 4) hybrid, a combination of the judgmental, empirical, and analytical methods.
Sabouhi et al. [17] proposed a framework for resilience assessment of power systems that provides a step-by-step process for assessing the Resilience Index (RI) using fragility curves. Panteli et al. [33] presented a similar framework to evaluate the resilience of a system using fragility curves and a sequential Monte Carlo Simulation method. Work in [31] introduced a cost-effective approach for system operators to calculate RI using fragility curves. One drawback of the detailed engineering approach is that it requires comprehensive information about the power network and its components, which is generally not publicly available data.
2.2 Data-driven Resilience Assessment
The data-driven resilience assessment approach, commonly known as the system-level approach, employs comprehensive data analytics to assess resilience metrics. It focuses on assessing the performance of the system that influences its ability to withstand and recover from disruptions. Operational resilience describes the capacity of a power system to sustain generation availability and continuous electricity delivery to consumers during and after disruptive events such as disasters, whereas infrastructure resilience focuses on the power system’s physical robustness and minimizing component failures or disruptions. Together, these aspects are reflected in probabilistic resilience curves to characterize system performance across a range of disturbance types, impacts, and severity [24],[8].
2.2.1 Resilience Triangle
The resilience triangle theory was developed by Bruneau et al. [7, 20] and serves as the foundation for resilience assessment. It primarily examines how a system’s performance deteriorates throughout an event. According to Tierney and Bruneau [7, 20], a system exhibits resilience when it minimizes the probabilities of failure and the impacts of failures, including damage, loss of life, and negative social and economic consequences, while also achieving shorter recovery times.
The resilience triangle method, depicted in Figure 3, provides a quantitative assessment of resilience via a time-dependent metric. The resilience metric evaluates system performance on a scale from 0% (a complete loss of service) to 100% (no disruption in service). The targeted performance is represented by $P_{T}$, while actual performance for operational and infrastructure resilience is denoted by $P_{op}(t)$ and $P_{in}(t)$ accordingly. As depicted in Figure 3, an extreme event occurring at time $t_{1}$ may cause substantial damage to the network, leading to a significant reduction in quality when it reaches $P_{op}^{min}$ (for operation) or $P_{in}^{min}$ (for infrastructure) at time $t_{2}$. After the event, recovery progresses until full restoration at $t_{4}$.
Figure 3: Schematic diagram of the resilience triangle.
A time-dependent resilience metric, $RI_{s1}$, is expressed in (3), which evaluates system resilience using both the resilience triangle and trapezoid. This metric is derived from the ratio of the area under the actual performance curve, $P_{op}(t)$, to the area between the target performance curve, $P_{T}$ [21, 87]. This approach can be applied to both operational and infrastructure resilience indexes,
$RI_{s1}=\frac{\int_{t_{1}}^{t_{4}}P_{op}(t)dt}{\int_{t_{1}}^{t_{4}}P_{T}dt}.$ (3)
In addition to the normalized metric presented in (3), Gholami et al. [22] introduced another metric,
$RI_{s2}$ $=\frac{\int_{t_{0}}^{t_{4}}[P_{T}-P_{op}(t)]dt}{\int_{t_{0}}^{T}P_{T}dt}=1-\frac{\int_{t_{0}}^{t_{4}}P_{op}(t)dt}{\int_{t_{0}}^{t_{4}}P_{T}dt}=1-RI_{s1}.$ (4)
In contrast to $RI_{s1}$, $RI_{s2}$ quantifies the reduction in resilience from the ideal value of 1 due to an extreme event, where zero signifies a fully resilient system and one represents a non-resilient system.
Moreover, G. Cimellaro et al. [23] articulated the resilience triangle metric for infrastructure to extreme events as
$RI_{s3}=\int_{t_{1}}^{T}\frac{P_{in}(t)}{T-t_{1}}dt,$ (5)
where
$P_{in}(t)=$ $[1-L(I,T_{RE})][H(t-t_{1})-H(t-(t_{1}+T_{RE}))]$
$\times f_{REC}(t,t_{1},T_{RE}),$ (6)
and
$T_{RE}=$ $t_{1}-t_{4}.$ (7)
Here, $L(I,T_{RE})$ represents the loss function, wherein $I$ and $T_{RE}$ denote the earthquake intensity and recovery time from the event, while $H(.)$ denotes the Heaviside step function. The parameter $T$ stands for the control time of the system after the restoration state. Additionally, $f_{REC}(t,t_{1},T_{RE})$ characterizes the recovery function. The authors emphasized metrics to measure resilience dimensions such as rapidity, redundancy, robustness, and resourcefulness.
While the resilience triangle is highly valued for its applicability across various systems, it lacks the capability to measure other pivotal resilience dimensions encountered by standard power systems. For instance, it fails to capture the duration infrastructure spends in the post-degradation state (from $t_{2}$ to $t_{3}$ in Figure 4) before restoration efforts begin and conclude [8]. To address these limitations, the resilience triangle concept has evolved into the resilience trapezoid framework [10], [13].
2.2.2 Resilience Trapezoid
The resilience trapezoid is the enhanced model of the resilience triangle, which incorporates the system’s degraded state in the event of delayed restoration actions, thereby providing a more realistic and comprehensive approach to understanding and measuring resilience [8],[25]. Illustrated in Figure 4, a conceptual resilience trapezoid delineates the states of a power system in response to an external disruption, alongside the temporal progression of these states over time [15], [25].
Figure 4: Schematic diagram of the resilience trapezoid.
Arjomandi-Nezhad et al. [26], proposed a recovery-based resilience metric which is defined as the ratio of the number of recovered customers to the average outage time of the affected customers
$RI_{s4}=\frac{P_{T}-P_{op}^{min}}{\frac{\int_{P_{op}^{min}}^{P_{T}}tdP}{P_{T}-P_{op}^{min}}},$ (8)
which can be rewritten as
$RI_{s4}=\frac{(P_{T}-P_{op}^{min})^{2}}{(T-t_{2})\times P_{T}-\int_{t_{3}}^{T}P(t)dt-\int_{t_{2}}^{t_{3}}P(t)dt}.$ (9)
In (9), $T$ represents the time at the end of the post-restoration state and $P_{op}^{min}$ represents the minimum level of the performance parameter during an extreme event. Similar to (8) and (9), equivalent formulations can be developed for assessing the resilience of system infrastructure. The model aims to restore as many loads as early as possible by maximizing the area under the recovery curve to achieve a higher resilience value.
Wang et al. [27] assessed $RI$ to enhance power resilience by finding the optimal size of Micro-grids (MGs), considering both cost and resilience factors. The authors introduced a version of $RI$ that denotes the proportion of weighted served loads after an extreme event, as expressed in (10). A smaller $RI$ indicates poor MG performance, and an $RI$ of 1 indicates no load shedding.
$RI_{s5}=\frac{\int_{t_{1}}^{t_{4}}P_{op}(t)dt}{P_{T}\times(t_{4}-t_{1})},$ (10)
where
$P_{T}(t)=w_{c}P_{c}+w_{n}P_{n},$ (11)
$P_{op}(t)=w_{c}P_{c}(t)+w_{n}P_{n}(t).$ (12)
Here, $P_{c}(t)$ and $P_{n}(t)$ indicate the power served to critical and non-critical loads. By assigning higher weights to critical loads, the metric shows that the system can remain resilient even if several non-critical loads are disconnected during the disturbance.
3 Resilience of Interdependent Power and Community Infrastructure Systems
Modern power systems operate within an interconnected ecosystem of critical infrastructures whose joint performance determines overall resilience. While power system resilience traditionally refers to the grid’s ability to withstand and rapidly recover from high-impact, low-probability (HILP) events while maintaining operational stability and infrastructure integrity [76], community functionality and broader infrastructure recovery are inherently linked to this capability. Community resilience—defined as the capacity of populations and institutions to prepare for, absorb, and recover from disasters—depends heavily on the continuity of electricity supply and the robustness of interdependent systems such as transportation, water, gas, and telecommunication networks [75]. In highly electrified and infrastructure-reliant societies, disruptions in the power grid can cascade across sectors, amplifying socioeconomic and operational impacts. Accordingly, assessing resilience now requires a system-of-systems (SoS) perspective that captures both technical inter-dependencies and community consequences. Table 1 summarizes representative studies examining the interdependence among the electric power systems and other communities and critical infrastructures.
3.1 Socioeconomic and Behavioral Dimensions
The community-level impact of outages varies widely based on demographic and socioeconomic conditions. Social vulnerability reflects the intersection of economic status, education, age structure, health, and access to critical resources, which collectively determine the capacity of communities to recover from disasters [102, 103]. In the context of power systems, social vulnerability determines how technical failures propagate into human and economic consequences: two communities experiencing identical outage durations may face vastly different recovery trajectories depending on their adaptive capacity and institutional support. This makes social vulnerability an essential bridge between grid performance and community resilience [75, 101].
A growing body of research has quantified how socioeconomic and behavioral factors shape outage exposure, duration, and subsequent recovery. A data-driven Social Vulnerability Index (SoVI) framework in [70] combined Principal Component Analysis (PCA), Adaptive Random Forests (ARF), and variance-based sensitivity analysis (VBSA) to identify the dominant societal drivers of outage vulnerability. Projected SoVI values for 2030 indicate a reduction in vulnerability disparities compared to 2022, reflecting long-term adaptation and incremental infrastructure improvement. Similar disparities were documented in [113], where a 0.1 increase in socioeconomic vulnerability was associated with a 6.1% increase in outage duration. Counties with the longest restoration delays (+6 hours above the mean) exhibited very high SoVI values ($\text{SoVI}\approx 0.91$).
Outage exposure also differs across racial and ethnic lines. A multilevel Bayesian analysis in [114] showed that counties with high racial/ethnic vulnerability $(\text{SoVI}>0.9$) experienced the longest outage durations during extreme weather events, including hurricanes, heat waves, and floods. Similarly, a spatially resolved analysis in [115] found that counties with high SoVI experienced approximately two times more outages lasting $\geq 1$ hour and approximately three times more outages lasting $\geq 8$ hours compared to low-SoVI counties.
Behavioral factors further amplify these disparities. A survey study in New York City [72] found that although 58% of respondents reported feeling generally prepared for outages, only 32% felt capable of enduring multi-day interruptions. Users of durable medical equipment (DME) were more prepared (70%), yet 74% still feared serious health complications during outages. These findings reveal a crucial gap between perceived and actual preparedness, especially for medically vulnerable or socially isolated residents. Additional household-level evidence from [116] shows that minority households reported significantly longer outages, while lower-income households experienced almost twice as many outage events. Table 1 summarizes the interdependency between power systems resilience and SoVI.
Age and health status further influence outage-related risk. Children, older adults, and medically fragile populations experience disproportionately severe impacts from prolonged outages and associated cascading failures [73, 71]. Documented public health risks include cardiovascular events, respiratory complications, renal failure, heat stress, gastrointestinal illness, and increased mental health burdens as outage durations rise [73]. To mitigate these outcomes, recent work highlights the importance of community-scale renewable energy resources and medically focused resilience planning to maintain essential electricity access during HILP events [71]. At the same time, emerging socio-technical models increasingly incorporate behavioral and human-response data, enabling AI-driven forecasting of consumer actions under uncertainty [79].
3.2 Infrastructure Interdependence and System-of-Systems Perspective
Modern power systems are part of a multilayer infrastructure network in which failures can cascade rapidly from one subsystem to another [101]. As illustrated in Figure 5, the electric grid is tightly coupled with telecommunications, natural gas and heating networks, banking and finance, transportation, water systems, emergency services, and government operations [77, 78]. Each subsystem retains partial operational independence yet depends on shared flows of energy, information, and logistics for sustained functionality. Under normal conditions, such coupling enhances system efficiency through coordinated control, information sharing, and optimized resource allocation. However, during HILP events, these same inter-dependencies accelerate disruption propagation and amplify cascading failures across sectors [80].
From a resilience perspective, these coupled infrastructures represent both stabilizing supports and potential pathways for systemic collapse. Telecommunications enable situational awareness and remote control; gas and water networks supply essential fuels and cooling resources for generation; and transportation systems facilitate repair crew mobilization, logistics, and power system recovery activities [100]. Consequently, resilience evaluation must adopt a SoS viewpoint that captures interdependence across sectors instead of assessing the power grid in isolation. Recent studies highlight this need by quantifying cross-sectoral sensitivity, recovery delays, and performance degradation under multiple hazards.
Water infrastructure exhibits strong dependence on electricity for pumping and pressure maintenance. Optimization-based resilience modeling in [117] showed that placing distributed generators (DGs) near pump stations can increase Water Distribution System (WDS) resilience by 13.35%, while local backup generation yields an additional 3.58% improvement. Under more severe disruptions, potable water systems experience drastic degradation when electrical service is lost. Using the Water Network Tool for Resilience (WNTR) with Pressure-Dependent Demand modeling, [119] demonstrated that hurricane-induced power outages reduce potable water system resilience by nearly 50%. These findings underline the critical reliance of water infrastructure on grid stability and the importance of backup power for maintaining essential public health services.
Interdependency between electricity, natural gas, and heating systems is equally impactful under extreme events. Dynamic optimal energy flow modeling in [118] showed that incorporating gas storage improves gas system resilience by 29.5% and that adding backup generators increases heating-system resilience by 15.6% following hurricanes. Multi-hazard Monte Carlo simulations in [120] further revealed that wind and earthquake events reduce outage energy resilience from 89% (wind) and 83% (earthquake) to 75%, accompanied by a 76%.
Transportation infrastructure also relies heavily on the power grid for operational continuity. Seismic resilience simulations in [75] showed that increasing earthquake magnitude from $M=6$ to $M=8$ reduces transportation network resilience by nearly 50%. Complementary robust optimization studies in [74] demonstrate that strategic hardening of critical power lines and placement of DGs significantly reduce disruption costs and preserve the operability of transportation-critical loads such as traffic signals, particularly during HILP events.
Linear regression analysis of post-hurricane recovery in [121] quantified strong cross-sector dependencies during restoration. Water, telecommunications, and hospital systems relied on power recovery for 89%, 93%, and 77% of their restoration trajectories, respectively. These findings affirm that power restoration serves as the primary driver of cross-infrastructure recovery and underscore the dominant role of electricity as the backbone of community resilience.
Figure 5: Interdependency of critical infrastructure systems as a SoS.
3.3 Resilience Hubs for Community Infrastructure Support
Resilience hubs provide a comprehensive, community-centered approach to addressing these multifaceted vulnerabilities. They recognize the interconnected nature of critical community infrastructure systems and the socioeconomic pathways through which HILP events affect community members. Resilience hubs also integrate public health and safety goals with technical considerations for power, transportation, communications, and essential services.
Resilience hubs are community-oriented facilities, such as schools, community centers, or libraries, that provide everyday public services during normal operating conditions and serve as disaster shelters, resource centers, and infrastructure support during emergencies (Fig. 6). During extreme events, they provide protected environments with physical safety, reliable power, thermal comfort, medical support, communications access, and other essential services [122].
Figure 6: Schematics of a Resilience Hub.
Beyond these community-facing roles, resilience hubs reflect the interdependence of community and grid resilience, addressing both immediate individual social vulnerability concerns and technical system-level performance issues [123]. A resilience hub may focus on supporting medically vulnerable populations, stabilizing public health services, ensuring access to cooling or clean water, or enhancing grid reliability using Distributed Energy Resource (DER) and adaptive load management. In doing so, resilience hubs directly address the intertwined technical and social factors that shape disaster impacts and recovery.
Resilience hubs are equipped with resilient backup power systems that not only ensure uninterrupted on-site access to critical services but can also provide grid-supporting functions [123]. This flexible capacity can stabilize nearby interdependent community infrastructure by supporting water pumping stations, emergency communications, transportation assets, and reliable electric vehicle (EV) charging during outages or evacuations. As a result, resilience hubs can be strategically located either to optimize accessibility for residents [124] or near critical infrastructure such as transportation corridors, water distribution systems, gas networks, and emergency services to help minimize cascading failures during grid disruptions [125, 123].
The framework in [124] designs and evaluates energy systems for resilience hubs in a region prone to extreme heat, identifying communities with limited access to air conditioning and heightened vulnerabilities to power outages. The resulting resilience hub design incorporates a DER portfolio sized to maintain safe indoor temperatures during a heatwave for up to four days without relying on the grid for power. Similarly, the grid-aware optimization in [126] demonstrates how hub-centered microgrids can mitigate outage impacts by balancing local resilience benefits with broader system constraints, demonstrating the role of hubs as both community assets and grid-supporting resources. These examples illustrate how resilience hubs bridge the gap between technical grid robustness and community resilience.
Table 1: Summary of the interconnection between community resilience and power systems resilience.
Category Community / Infrastructure Methodology Key Findings Extreme Event
Behavioral[70, 72, 113, 114, 115, 116] Older adults, Low-income households, DME users Telephone survey [72] 58% reported outage preparedness; 70% of DME users were prepared, but 74% had serious health concerns during outages General climate change
Home EV chargers, Educational attainments, Low income residents SoVI analysis using PCA, ARF, VBSA, Sobol [70] Most influential factors affecting outage vulnerability: Education (%99.7) — higher levels decrease vulnerability; Affected Customers (%86.8) — larger outage exposure increases vulnerability Heatwaves, storms
Low-income, elderly, socially vulnerable SDM regression using SoVI [113] 0.1 increase in socioeconomic vulnerability results in a 6.1% increase in outage duration; Counties with the highest delays in power restoration—+6 hours longer on average—had very high SoVI = 0.91 Hurricane
High racial/ethnic vulnerability Multilevel Bayesian modeling using gamma regression [114] Counties with SoVI >0.9>0.9 faced the longest outage duration Hurricanes, heat waves, floods, winter storms
Urban centers, rural counties SoVI analysis using spatial cluster, and non-parametric statistical tests[115] High-SoVI counties experienced approximately $2\times$ more outages lasting $\geq 1$ hour and $3\times$ more outages lasting $\geq 8$ hours compared to low-SoVI counties General climate change
Diverse urban households Household surveys using Kruskal–Wallis tests[116] Minority households reported longer outages: 58.1% of minority respondents compared to 40.4% of white respondents; Low-income households experienced more frequent outages, with 17.2% versus 8.4% among higher-income households General climate change
Infrastructure[117, 118, 119, 75, 120, 121] Water distribution systems Resilience metrics computation using SOCP optimization [117] Placing DGs near pumps increased RA by approximately 13.35%, indicating stronger WDS performance relative to PDS degradation; Enabling local backup generation increased RA by about 3.58%, indicating a modest improvement in maintaining WDS performance Storm
Natural Gas, and Heating System Dynamic optimal energy flow modeling with resilience metrics[118] Gas system resilience improved by 29.5%— when gas storage capacity was considered; Heat system resilience increased by 15.6%—after adding backup generators Hurricane
Potable water system WNTR tool with Pressure-Dependent Demand [119] Power outages reduced potable water system resilience by 50% (e.g., water service dropped below 50% within 8.8 days), and extended full recovery time by over 300% (e.g., from 6.7 to 28+ days) Hurricane
Urban transportation (road network) Seismic resilience using Monte Carlo [75] Resilience drops sharply as earthquake magnitude increases — at $M=8$, network resilience is approximately 50% of its value at $M=6$, and recovery times are about 15 times longer Earthquake
Natural gas, and Heating system Monte Carlo simulation of component failures[120] Total energy resilience fell from 89% (wind), 83% (earthquake), to 75% (combined) — a 16% drop under multi-hazards; Energy curtailment reached 76% under multi-hazard conditions and required 52% longer restoration time Wind Storm, Earthquake
Water system, Telecommunications (cell sites), and Hospitals Linear regression analysis of recovery curves[121] Water, telecommunication, and hospital systems rely on power recovery for 89%, 93%, and 77% of their recovery, meaning only 11%, 7%, and 23% of their resilience is independent of the power grid Hurricane
4 Power Systems Resilience Enhancement Strategies
Resilience assessment provides quantitative methods to identify vulnerabilities and strengths of power systems, enabling proactive strategies to strengthen resilience before extreme events occur. Here, different types of resilience-enhancing strategies are identified and explained.
4.1 Hardening Power Network
System hardening refers to modifications made to power system infrastructure to minimize vulnerability to severe events [16, 12]. Hardening is particularly effective at mitigating the impact of extreme events in distribution systems. Typical hardening techniques include reinforcing substations, upgrading or undergrounding power lines, and armoring poles [16, 12, 30]. The U.S. National Electric Safety Code requires the replacement of poles when their structural ability falls below two-thirds of their original strength [30].
In [18, 39], a tri-level optimization is proposed that focuses on hardening power networks to enhance the resilience of distribution systems. It explores strategic decision-making for system hardening through scenario planning under maximum-damage conditions. System hardening, while serving as an effective resilience enhancement strategy, involves a significant financial burden [16, 12]. Consequently, substantial research has focused on developing cost-effective hardening solutions without compromising safety or efficiency [12]. The integration of DER into modern power systems introduces additional challenges. The uncertainty associated with renewable energy operation complicates traditional scenario-based hardening techniques, as it increases the number and severity of system contingencies and potential worst-case scenarios [30].
4.2 Resource Allocation
Resource allocation is a resilience enhancement approach that involves strategically deploying DG units, mobile power sources, and other critical resources at key nodes in the power system to bolster resilience [42]. Given the uncertainty of DER operation during extreme weather conditions, it is essential to back up the lost power to maintain system stability. This helps strengthen grid resilience while maintaining reliable operation during both normal and extreme conditions [40]. Strategically positioning DERs within the network can reduce the need for critical load curtailment and enhance overall system resilience [42, 43]. In [41], a two-stage stochastic mixed-integer programming model was developed to optimize the location and sizing of DGs (in the first stage), minimize the operating costs of DGs and load shedding (in the second stage). Other design decisions, like optimal DG size, can be made using similar optimizations with the objective to maintain resilience above a specified threshold [27].
4.3 Optimal Scheduling of Repair Crews
After an extreme incident, quick repairs of the power system are needed to minimize load-shedding losses. Maintaining resilience requires a proactive repair crew deployment approach to prepare for post-disaster repair and restoration. Optimization models have been proposed to minimize load shedding by sending repair crews to the ideal locations before extreme events. [44]. Their goal is to maximize the expected restored load while minimizing the repair duration. In [45], Anmar Arif et al. propose co-optimization of repairs, reconfiguration, and DG dispatch in a two-stage method for outage management.
4.4 Network Reconfiguration
Network reconfiguration refers to the modification of the network topology by opening normally closed sectionalizing switches and closing normally open tie switches. In order to follow a set of operation constraints, network reconfiguration problems are typically modeled using various objectives, like minimum switching operations, operation costs, or power loss, and are subject to load balancing and network constraints [28]. A number of technologies are employed to improve reconfiguration capacity and distribution system resilience, including smart transformers, on-site DG units, integration of DERs, fault prevention devices, and fault detection systems [12]. Network reconfiguration can be particularly effective at improving power system resilience when more sophisticated measuring instruments are used. Devices like distributed remote controllers, phasor measurement units (PMUs), communicating relays, reclosers, advanced supervisory control and data acquisition (SCADA), and outage management systems (OMS) allow for remote reconfiguration, which can reduce load interruption times. A more effective approach uses the real-time status of the system during each stage of an event to guide subsequent decisions [30].
5 AI Methods for Power Systems Resilience
AI and ML methods are emerging as powerful power system operation and planning tools applied to a variety of tasks such as load forecasting, outage prediction, renewable generation prediction, demand response management [47], anomaly detection [89], stability assessment, control, and restoration [46]. Most ML techniques can be categorized into four groups: 1) Traditional ML, 2) Artificial Neural Network (ANN), 3) Reinforcement Learning (RL), and 4) Advanced Generative and Language Models. Table 2 summarizes different machine learning approaches for power system resilience, highlighting their performance, limitations, and related extreme events.
Power outage prediction has emerged as a critical area of research to bolster the resilience of energy systems against various disruptions, including weather events, equipment malfunctions, and grid instabilities. Logistic regression is a simple, fast, and robust method for outage prediction [51].
Some techniques offer higher accuracy than logistic regression. For instance, Support Vector Machines (SVM) has been utilized to predict outages by assessing the condition of grid components, with SVM generally outperforming logistic regression in terms of accuracy [50, 51]. In [50], a decentralized framework is proposed to isolate affected regions during hurricanes and wildfires. One other notable method is to use ensemble learning techniques. For example, ADABOOST+ has been used to analyze the impacts of lightning and wind on power outages across four cities in Kansas [49]. ADABOOST+ has outperformed ADABOOST.RT, traditional regressional models [86, 51], and neural networks [85].
Convolutional Neural Networks (CNN) combined with long-short-term memory (LSTM) have been applied to predict outages across selected counties in New York, studying the socioeconomic impacts on vulnerable populations based on household income and racial demographics [52].
In [48], the authors enhance prediction model accuracy by pre-classifying events into three severity levels using Quantile Weight Distance (QWD): low (<< 100 outages), moderate (100–1,000 outages), and high (>> 1,000 outages). A conditioned outage prediction model (OPM) is then developed using a Bayesian Additive Regression Tree (BART), Random Forest (RF), and Ensemble Decision Tree (ENS) to outperforms the standard OPM. Addressing data preprocessing challenges, a comprehensive approach for data collection and feature selection is proposed, which includes environmental features, infrastructure specifics, drought, soil moisture, pressure, temperature–humidity, precipitation, convection, and wind [53]. The study suggests that the combination of log-transform, synthetic minority over-sampling technique (SMOTE), and conditional bias correction (ModelF) provides higher accuracy (above 80%). The role of socioeconomic variables in power outage prediction is also examined [88].
Table 2: Summary of machine learning in different studies to enhance power resilience
Category Methodology Key Findings Drawbacks Extreme Event
Traditional ML Regression-based Model [86, 51, 90, 91] Simple or linear regression accuracy typically 60–70%, where higher-order and weighted logistic regression can exceed 90%. Event-based (customer affected): 87-93% [91, 51], hourly (binary classification of outage probability): 60-75% [90], daily (outage counts): 60-75% [86] Tend to underestimate high-impact low probability events Wind and lightning events [86], Hurricanes (simulated) [51], Hurricane Ian (Florida, 2022)[91], Extreme weather outages [90]
Ensemble Learning [49, 92, 93] Predictive accuracy ranging between 65–99%; boosted ensemble achieves higher accuracy by adaptive weighting. Daily (outage counts): 65-80% [49, 92], event-based (binary classification): 95-99% [93]. Model lacks explicit representation of network topology Wind and lightning events [49], Hurricanes and storms (North Carolina)[92]
Artificial Neural Network (ANN) Classical ANN [85, 94] High predictive accuracy (>>80%), outperforming regression due to nonlinear mapping ability. Daily (outage count): $\approx 80\%$, seasonal (outage count): 99%. [94]. Prone to overfitting with limited data Wind and lightning (Kansas)[85], Rainfall, storms [94]
Deep learning (CNN, LSTM, GNN) [52, 65, 88] Achieved 90–99% accuracy; adding additional hidden layer improves model accuracy. Hourly (customer affected): 91-95% [52, 88], Event-based (binary classification): 92-99% [14]. Requires large, diverse datasets Hurricanes, floods (New York)[52], High winds and Ice-storms (Michigan)[88]
Reinforcement Learning (RL) Classical RL [62, 60, 57] Effective for system restoration, control, and decision making, serving as an action-based framework that emphasizes optimal decision, however, its application to outage prediction remains limited. Use small-case network Hurricanes, Storms [60, 62, 57]
Deep RL [38, 55, 61, 54] Integrates neural networks to learn optimal actions in large networks where classical RL struggles with high-dimensional states. Resilience indices increase by 5–15% and reduce computation time. Computationally expensive and less interpretable Natural disaster[55, 61], Cyber-attack [54], Hurricane [38]
Advanced Generative and Language Models Transformer-based [99, 96] Reduce forecast errors compared with numerical models. Event-based (outage count): 84-99% [99]. Self-attention mechanism is limited Natural disaster, cyber-attack.
LLM-based Models [97, 98, 96] Extends transformer-based model, achieve 90–99% accuracy through large-scale parameter tuning and contextual adaptability. Event-based (customer affected): (>>90%) [97, 98]. High computational and storage cost Hurricanes, cyber-attacks, natural disasters.
RL has emerged as a prominent technique in stability control and restoration tasks. In RL applications, the environment is typically modeled as a finite-horizon Markov decision process, defined by the tuple $(S,A,p,r,d_{0},\gamma)$. Here, $S$ and $A$ represent the state and action spaces, respectively. $p(s^{\prime}\mid s,a)$ is the state transition function indicating the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$; $r(s,a)$ is the reward function for taking action $a$ in state $s$; $\gamma$ is the discount factor for future rewards, and $d_{0}$ is the initial state distribution. The results demonstrate that deep reinforcement learning (DRL) outperforms the greedy algorithm and can control voltage regulation [54].
Different RL tasks are characterized by unique state and action spaces, along with task-specific reward functions. [55, 56, 57] focused on using RL for MG formation post-disaster, whereas [58, 59] further explore MG operation. These models often have discrete action spaces (e.g., toggling switches on and off for specific DERs), and sometimes continuous spaces (e.g., adjusting setpoints for selected DERs) [57, 54].
RL has been applied to optimize restoration strategies, such as scheduling repairs for components [60], dispatching mobile energy resources [61, 62], and managing repair crews [62]. [60] compares the performance of the Q-learning algorithm with Greedy Search (GS) and Exhaustive Search (ES). Q-learning outperforms GS and ES by achieving optimality faster.
Extreme events can disrupt communication infrastructure, making centralized decision-making impractical. To address this, a decentralized framework is proposed in [62] for deploying DERs. The proposed method outperforms other models in the resilience index by 6-9%. However, it does not account for uncertainties in DERs, and its optimal performance remains 3.53% lower than mixed-integer linear programming (MILP).
Beyond RL, other ML approaches are also instrumental in enhancing power system resilience. Predictive control policies for MG management have employed tree-based models to forecast load demands and solar production [63]. Bayesian Networks and hybrid statistical algorithms like the Fragility-curve Monte Carlo Simulation Scenario Reduction (FC-MC-SCENRED) have been compared for pre-hurricane system optimization for distribution networks [64]. Graph neural networks (GNNs) have been proposed to assess node criticality and identify vulnerable lines in cascading outage studies [65, 66]. For disturbance management, CNNs paired with autoencoders have been developed to dynamically adjust shunt capacitors and transformers during electromagnetic pulse disturbances [67]. A one-dimensional CNN has been designed to detect data intrusions in smart meters [68]. Advanced deep learning techniques have been applied for system-wide online load restoration with high wind power penetration [55], and MG planning in active distribution systems using Bayesian Regularization Backpropagation [69].
Recent advancements in transformer-based generative intelligence offer immense potential to apply in the power systems to better capture complex patterns between parameters. This transformer-based model can tune millions of parameters at a time, which provides the ability to represent more accurate results [96]. The models such as eGridGPT and Weather Foundation Model, are used to train and forecast more accurate power grid analysis and decision making [95]. Domain-specific large language model (LLM) framework extends beyond conventional predictive learning by introducing reasoning, context awareness, and multi-modal understanding of the grid’s characteristics [96]. Transformer-based generative architectures demonstrate that self-attention mechanisms can effectively combine information from different sources and understand how events change over time and locations [99, 98]. However, the model requires high-performance computing infrastructure to process datasets for training and real-time inference. In addition, the model may provide inaccurate output and need verification. Despite these challenges, these models offer a significant opportunity to deliver more accurate and adaptive analyses of power systems behavior under various operating conditions.
6 Techno-Legal Aspects of Resilient Power Systems and Communities
The enhancement of power systems resilience, as reviewed in this paper, transcends overwhelmingly technical implementation. This focus is critical, as power systems function as the keystone infrastructure within a complex SoS. This dependency elevates the power grid to a primary technical and regulatory priority, as a failure cascade originating from the grid can paralyze an entire community. This section, therefore, bridges the resilience assessment methods, AI applications, and interdependency models discussed previously with the evolving transatlantic legal regimes. We analyze the implications for system operators, municipalities, and community energy groups, and align these legal considerations with the paper’s focus. Figure 7 conceptualizes the power grid as the central keystone within a SoS architecture. The diagram illustrates how the techno-legal frameworks (e.g.North American Electric Reliability Corporation Critical Infrastructure Protection -NERC CIP- [104], NIS2 Directive [105]) impose a top-down regulatory structure (Tier 1) on the power system (Tier 2). This governance is a direct consequence of the bottom-up dependency (Tier 3), where all other critical societal infrastructures (healthcare, water, transport) fundamentally rely on the power system for their own operation.
Figure 7: Techno-legal framework for critical infrastructures as a part of SoS.
6.1 The Evolving EU Regulatory Landscape
The relevant European Union (EU) regulatory landscape is multifaceted. The GDPR establishes a strict regime for processing personal data, which is pertinent to outage analytics, Advanced Metering Infrastructure (AMI), and the use of social vulnerability inputs [106]. Concurrently, the NIS2 Directive imposes comprehensive risk management, supply chain security, and incident reporting obligations on essential entities within the energy sector [105]. This is complemented by the CER Directive, which mandates risk assessments and resilience measures to protect critical infrastructure. Finally, the forthcoming EU AI Act introduces a risk-based framework for AI systems, classifying safety components in critical infrastructure and AI models that affect grid services as high-risk, thereby imposing stringent duties regarding data quality, risk management, transparency, and human oversight.
6.2 Comparative US Regulatory Context
In the United States, a parallel framework exists, though it is often more sectoral. Mandatory cybersecurity standards for the bulk electric system are set by the NERC Critical Infrastructure Protection (CIP) standards, enforced by Federal Electric Regulatory Console (FERC) [104]. Broader critical infrastructure guidance and risk management frameworks are promulgated by the Cybersecurity and Infrastructure Security Agency (CISA) and the National Institute of Standards and Technology (NIST), particularly the widely adopted NIST Cybersecurity Framework (CSF)[107]. For data privacy, the US lacks a single federal equivalent to GDPR, relying instead on a patchwork of state-level laws (e.g. the California Consumer Privacy Act - CCPA) [108] and federal sectoral laws (e.g. Health Insurance Portability and Accountability Act of 1996, which becomes relevant when correlating outages with vulnerable health [109]). For AI governance, the US is currently focused on non-binding frameworks, notably the NIST AI Risk Management Framework (AI RMF) [110], which encourages a voluntary, evidence-based approach to managing AI risks, contrasting with the EU’s binding, compliance-driven AI Act [112].
6.3 Implications for Data and Models in Resilience Analytics
These legal frameworks have direct implications for the data and models utilized in resilience analytics. The processing of social and outage data, including health, demographic, or social vulnerability features, demands strict purpose limitation, aggregation, or the application of differential privacy under GDPR[106], or adherence to de-identification standards under HIPAA [109]. Privacy-preserving learning methodologies are preferable when correlating outage data with vulnerable populations, directly supporting the community-centric metrics reviewed. Furthermore, asset and interdependency data, such as grid topology, substation layouts, and linkages to gas, water, or transport systems, must be treated as sensitive operational data under NERC CIP [104] and NIS2[105]. Where real topologies are restricted for security reasons, validated synthetic networks should be employed to support the fragility-based and data-driven assessments.
6.4 Cybersecurity by Design for Resilient Operation
Achieving resilience necessitates a cybersecurity-by-design approach, integrating controls from frameworks like NIS2 [105] and NERC CIP [104] directly into resilience functions. Technical workflows for defensive islanding, DER dispatch, and network reconfiguration must be mapped to specific controls such as robust authentication, least privilege access, network segmentation, and secure update protocols. This extends to the secure supply chain for operational technology, including IEDs and PMUs, where requirements for Software Bills of Materials (SBOMs) [111] and vulnerability disclosure are essential. Incident response procedures must also be designed for cyber-physical scenarios, particularly those that degrade communications during critical operations like MG formation.
6.5 Law-Tech Integration in Interdependent Infrastructures
The techno-legal framework must directly address the SoS realities of interdependent infrastructures. Resilience planning should include joint exercises for power, gas, water, and health services, supported by data sharing templates that pre-define lawful bases (under GDPR) and confidentiality clauses, reflecting the cascading risks detailed in this review. In procurement and liability, tenders for grid equipment must embed relevant clauses (e.g., NIS2 [105], NERC CIP [104]conformity, AI Act-ready documentation [112]), requiring evidence of secure-by-design principles. For AI-assisted decisions, liability allocation must be supported by immutable logs linking sensor inputs, model outputs, and human approvals.
To ensure accountability, techno-legal performance indicators—such as time to detect and report cyber incidents or percentage of datasets with completed Data Protection Impact Assessments (DPIAs), should be tracked with the same rigor as traditional resilience indices [127].
7 Research Gaps, and Future Research Directions
Despite the development of various metrics to evaluate power system resilience and subsequent enhancement strategies, current methodologies have the following limitations.
  1. Comprehensive Failure Models: Current models like fragility curves do not adequately capture the spatiotemporal dynamics of natural disasters or their impact on grid components. Improved fragility curves, alongside scenario-based simulations, could lead to more accurate failure models [11].
  2. Data Utilization and AI/ML Algorithms: The advancement of AI has significantly enhanced the ability to process large datasets for forecasting outages and assessing resilience. However, the scarcity of historical data and the challenges of data collection remain significant barriers. There is a crucial need for collaboration between utilities and researchers to overcome these obstacles and for effective data integration and cleansing practices to utilize big data fully [29].
  3. Confidentiality and Data Sharing: The confidentiality surrounding vital infrastructure data often limits the availability for research. The real topology of power networks is often proprietary and highly sensitive. However, synthetic networks can be generated using publicly available datasets, which share similar properties with real network topologies [36].
  4. Extreme Weather Events Modeling: Due to the rarity of extreme weather events, it is difficult to collect adequately large extreme weather event datasets. With generative AI, extreme weather scenario data can be generated, which shares a similar distribution to the real extreme weather events [37]. Thus, the resilience of the power systems can be simulated before such events even happen in the real world. Planning for the recurrence of extreme events shortly after an initial event is identified as a key priority for future research.
  5. Integration of AL/ML with Stochastic Methods: Combining AL and ML techniques paired with stochastic approaches could significantly improve the modeling of nonlinear behaviors in power systems. This integration promises more precise predictions and assessments, particularly beneficial for complex and large-scale systems.
8 Conclusion
This paper offers a comprehensive review of community-centric power system resilience assessment, summarizing frameworks, characteristics, and enhancements. Resilience assessment and enhancement play a vital role in facilitating resilient planning and operation. The paper identifies two fundamental approaches to evaluating power system resilience. The first is detailed engineering: this approach focuses on developing the fragility curves of the components within a system to assess resilience. The second is data-driven: this method relies on aggregated outage data to evaluate resilience at a broader level, using frameworks such as the resilience triangle, resilience trapezoid, or resilience curves. The community-centric resilience assessment necessitates considering socioeconomic and demographic factors, as well as the interdependence between power and other energy systems. Next, the paper presents a comprehensive review of numerous research papers concerning the improvement of resilience within power systems. Moreover, the application of AI and ML technologies in power system resilience is explored, which facilitates the advanced analysis of system responses to extreme events.
Beyond technical metrics, this review shows that power systems sit at the center of a techno legal system of systems, where frameworks such as NERC CIP, NIS2, GDPR, and the EU AI Act shape how data, AI tools, and digital twins support resilience. Embedding these obligations into grid design, operation, and AI-driven decision support strengthens community-centric resilience while aligning outage analytics, interdependency modeling, and cyber-physical protection with legal accountability. Future work should integrate resilience indices with compliance-oriented indicators, linking system performance to measurable duties for operators and for other critical infrastructures that depend on the grid.